# A Survey on Monocular Re-Localization: From the Perspective of Scene Map Representation

### v1

@@ 7.3 Neural Radiance Field
NeRF implicitly represents a scene as a “radiance field”: a volumetric density that models the geometry of the scene, and a view-dependent color that models the appearance of occupied regions of the scene NeRF . A Multiple Layer Perception (MLP) is utilized to takes the 3D position of a point $\textbf{p}^{\mathcal{G}}=(x,y,z)$ and the unit-norm viewing direction of that point $\textbf{v}=(d_{x},d_{y},d_{z})$ as input, and estimate the density $\sigma$ and color $c$ of the point, $(\sigma,c)\leftarrow F_{\mbox{NeRF}}(\textbf{p}^{\mathcal{G}},\textbf{v})$. Scene map is represented by the parameters $\Theta$ of $F_{\mbox{NeRF}}()$. Given a set of RGB images $\{\mathcal{I}_{t}\}$ with known camera poses $\{\textbf{x}^{\mathcal{G}}_{t}\}$, NeRF is trained by minimizing a photometric loss $\mathcal{L}$ NeRF :
@@
NeRF learns to synthesize novel views associated with any camera poses, enabling virtual image rendering and scene geometry recovery, thus have been applied in MRL in recent years. Some works directly uses pre-trained NeRF to estimate pose while other works utilize NeRF as a novel data augmentation method to enrich the training data.
NeRF as pose estimator: As the first framework to estimate pose using NeRF, iNeRF [28] directly optimizes 6 DoF poses with freezed network parameters $\Theta$ by minimizing the photometric error between rendered pixels and observed pixels, which is the inverse problem of original NeRF NeRF :
@@
Such a joint pose estimation and NeRF training scheme then motivates pose-free NeRF Meng2021GNeRFGN ; NoPe-NeRF . In LATITUDE LATITUDE , authors trained a pose regressor through images generated from pre-trained NeRFs, which provides an initial value for hierarchical localization as VPR methods without the requirement of saving reference images. In fine pose optimization stage, LATITUDE LATITUDE minimizes the photometric error between the observed image and rendered image by optimizing the pose on the tangent plane. In Loc-NeRF LocNeRF , a pre-trained NeRF model is integrated into a Monte Carlo localization to assigns weights to particles by image similarity between query image and rendered image. In NeRF-Loc liu2023nerfloc , a NeRF conditioned on reference images takes 3D point as input and generates corresponding 3D descriptors so that 2D-3D matches can be obtained by matching. Pose is then estimated by PnP solver.
NeRF as data augmentation: As a more straightforward way to apply NeRF in MRL, NeRF can be used to render virtual images at unseen viewpoints so that the overfitting problem of APR methods can be effectively solved and thus improves performances of APR Moreau2021LENSLE ; DPN ; DFNet . LENS Moreau2021LENSLE applies pre-trained NeRF-W NeRF-W as an offline data augmentation pipeline to enhance APR model. As an online augmentation way, Direct-PoseNet DPN minimizes the photometric errors between query image and virtual image rendered at estimated pose while DFNet DFNet uses feature map errors DFNet during online localization at unseen viewpoints, making APR method more precise and scalable.
Summary: As a newly emerged technique, NeRF has two favourable characterises: 1) the implicit representation of scene; 2) the ability to render synthesis image at novel viewpoint. These two characterises present great potential in MRL researches. Current attempts in this area focus on incorporating NeRF with APR methods. With the assistance of NeRF, APR could perform comparable or even better than VL-MRL methods in small- and medium-scale scenes. But we have to say the utilization of NeRF in MRL is still an opening question to be explored.
@@
Figure 16: Instances of APR method enhanced by NeRF, a) Direct-PoseNet DPN using photometric consistency between query image and rendered image, b) DFNet DFNet using feature consistency between query image and rendered image.
8 Benchmarking Monocular Re-Localization
@@
In this section, the common evaluation metric and dataset are introduced, and then we provide the performances of some SOTA MRL methods with various type of scene map on these benchmark. Looking back to all kind of MRL methods mentioned above, some of MRL methods retrieve reference image as query pose approximation while others perform precise pose estimation for query image, as listed in Tab. 1. These two kinds of MRL are generally tested using different evaluation metric. We provide introduction and evaluation respectively.
8.1 Evaluation metrics
8.1.1 MRL  for  IPA
@@
Those retrieval-based MRL utilize evaluation metric as image retrieval task. A retrieved reference image/point cloud clip is defined as true positive if it was geometrically near the query images. The ground-truth is generally labeled by geometric distance based on GPS signals (for example, a threshold of 10 m in Li2023I2PRecRI ) or manually defined yue1 ; yue2 . With the ground-truth, we can measure Precision and Recall metric, where precision denotes the ratio between true positives and all the retrieved references and recall is the ratio between the number of true positives and all the potentially matched query-references existed in this scene. In some MRL application need to retrieve comprehensively, we measure the ratio of retrieved true positives over top-K candidates, i.e., Recall@K. In some applications that a mismatch will make the whole system dramatically fails, e.g., loop closure detection in SLAM LCD2022TITS , we turn to measure the maximum recall metric with 100% precision. The higher these two metrics, the better MRL performs.
8.1.2 MRL  for  FPE
@@
Direct measurement about estimated poses is applied in those MRL methods aiming at precise pose estimation. The common evaluation metrics are position error for translation $\textbf{t}^{\mathcal{G}}_{v}$ (in distance) and orientation error for rotation $\textbf{R}^{\mathcal{G}}_{v}$ (in degree) between estimated results and ground-truth poses. The lower the errors, the better MRL performs.
@@
Table 1: The categorisation of various MRL methods based on the function.
Function
Method
Initial Pose Approximation
VPR (Sec. 3.1) and I2P-VPR (Sec. 5.2)
Fine Pose Estimation
RPR (Sec. 3.2), VL-MRL (Sec. 4), G-I2P (Sec. 5.1), I2P-RPR (Sec. 5.2), I2P-MLoc (Sec. 5.2), MRL with HD map (Sec. 6), APR (Sec. 7.1), SCR (Sec. 7.2), and NeRF-based MRL (Sec. 7.3)
8.2 Dataset
@@
There are a large number of datasets available for the evaluation of MRL methods and it is difficult to give a complete overview. Here we only present some typical datasets that we have used to evaluate MRL methods in this survey.
8.2.1 MRL  for  VPR  only
@@
The dataset used to evaluate VPR is required to include query and reference images that can be paired by geometric distance. So the dataset should have traverse the same scene multiple times, or the trajectory of the dataset should be “looped”.
Gardens Point dataset GardenPoint2 has been recorded on the Gardens Point Campus in Queensland University of Technology at different times of day and night. It contains three traverses of scene that two in the day while one in the night.
SPEDTest dataset SPED includes frames depicting various locations from all over the world. The collected scenes are with a vast range of weather, seasonal, and illumination conditions.
Nordland dataset nordland was first introduced to VPR evaluation by seqslam . It represents a 728 km of train journey in Norway during summer and winter seasons. Nordland dataset represents natural (non-urban), outdoor environment, which is unexplored in any other dataset.
Pittsburgh dataset pittsburgh has two versions of test dataset for VPR evaluation. Pitts250k-test contains 8k queries and 83k reference images, collected from Google Street View and Pitts30k-test comprises 8k queries and 8k references. Both Pittsburgh datasets show significant viewpoint changes.
Mapillary Street-Level Sequences (MSLS) dataset MSLS has been collected using camera mounted on ground vehicles and presents a wide range of viewpoint and illumination changes.
Oxford dataset fabmap , including CityCentre and NewCollege, is collected through the fish-eye binocular images of a wheeled robot in a campus. Ground-truth matches between query and reference image are provided by the authors.
St. Lucia dataset RatSLAM is recorded through a suburb in Brisbane at 5 different times during a day, and also on different days over a period of two weeks. The viewpoint changes are limited, but significant appearance changes due to the different times of day are included, as well as some dynamic objects such as traffic or cars parked on the street.
17-Places dataset 17places consists of a number of different indoor scenes, ranging from office to labs, hallways, seminar rooms, bedrooms, etc. This dataset exhibits both viewpoint- and conditional-variations.
Except to these datasets, tremendous challenging datasets aiming at evaluating VPR still exists, such as Tokyo 24/7 dataset tokyo247 , ESSEX3IN1 dataset ESSEX3IN1 , INRIA Holidays dataset INRIA , Synthia dataset SYNTHIA , etc.
8.2.2 MRL  for  IPA
To evaluate both VPR and I2P-VPR, it requires the dataset to have both images and LiDAR sweeps. Since I2P-VPR is a newly emerged research, there is very few dataset developed for the evaluation in this field.
@@ 8.2.2 MRL  for  IPA
KITTI dataset KITTI is a widely used AD benchmark that contains 22 sequences, of which 11 sequences (00-10) present ground-truth trajectories that is originally used for VO or SLAM evaluation. Based on the given pose of each image, ground-truth matches between query and reference image/point cloud can be labelled, and both VPR and I2P-VPR can be evaluated since synchronized LiDAR scans are also provided, we can offline build LiDAR point cloud map using LiDAR SLAM to generate reference point cloud clip as mentioned in Sec. 5.2.
8.2.3 MRL  for  FPE
Some datasets provide ground-truth pose for each image so that we can evaluate the MRL methods aiming at FPE. These dataset usually provide full sensor suite, so we can build visual landmark map (Sec. 4), point cloud map (Sec. 5), HD map (Sec. 6), and even DNN-based implicit map (Sec. 7).
Aachen Day-Night benchmark6dof utilizes reference images to build the scene map, which were taken during daytime with hand-held cameras over a period of about two years. The datasets offers query images taken at daytime and at nighttime so that viewpoint changes and illumination changes are both included. It is popularly used in evaluation for VL-MRL methods.
InLoc Dataset InLoc is designed for large-scale indoor localization that contains significant variation in appearance between queries and the 3D database due to large viewpoint changes, moving furniture, occlusions or changing illumination. The dataset is composed of a database of RGBD images geometrically registered to the floor maps augmented with a separate set of RGB query images taken by hand-held devices to make it suitable for the task of indoor localization. The reference images are 9,972 perspective images consisting of 277 RGBD panoramic images obtained from scanning two buildings at the Washington University in St. Louis. The query images consist of 329 photos using a smart-phone camera annotated with ground-truth 6 DoF camera poses. It is used in evaluation for VL-MRL methods.
KITTI Dataset KITTI provides ground-truth pose for each image, so it can also be used to evaluate MRL for FPE. Since multiple sensor data is provided, it can support evaluation for many kind of MRL.
KAIST Urban KAIST provides LiDAR data and stereo image with various sensors targeting a highly complex urban environment. It captures features in urban environments (e.g. metropolis areas, complex buildings and residential areas). Wen et al. build HD map for it so that it can be used to evaluate HD-MRL methods TM3Loc .
7Scenes posenet is a collection of tracked RGB-D camera frames. The authors used KinectFusion system newcombe2011kinectfusion to obtain the ground truth camera poses, and a dense 3D model. Several sequences were recorded per scene by different users, and split into distinct training and testing sequence sets. It supports the evaluation of almost all the kinds of MRL methods.
Cambridge Landmarks posenet is a large scale outdoor MRL dataset taken around Cambridge University. It contains original video, with extracted image frames labelled with their 6 DOF camera pose and a visual reconstruction of the scene. Similar to 7Scenes posenet , it also supports the evaluation of almost all the kinds of MRL methods.
The mentioned above datasets only cover a small part of dataset for evaluating MRL methods for FPE. Except to these dataset, the famous dataset, CMU Seasons and Extended CMU Seasons benchmark6dof , RobotCar Seasons benchmark6dof , ETH-Microsoft Dataset eth_ms_visloc_2021 is also widely used in evaluation, and Map-free Localization dataset arnold2022mapfree provides a specialised benchmark for evaluating RPR methods when only one reference image is included in the scene map.
8.3 Performance of SOTAs
8.3.1 MRL  for  IPA
MRL methods for IPA, including VPR and I2P-VPR, tend to retrieve reference images given query image, so they are measured and compared based on the precision and recall metrics of their retrieved results.
VPR: First, we give some performance of VPR methods in public datasets, evaluated with different metrics, e.g., recall@K (in Tab. 2) and maximum recall at 100% precision (in Tab. 3). It shows in Tab. 2 that the differentiable VLAD pooling layer could effectively aggregate dense local descriptors into global features, thus making NetVLAD and its extensions SPE-NetVLAD ; vlaad perform better than average pooling netvlad and GeM pooling GeM in retrieval-based GF-VPR tasks. Then, benefit from the global fusion of features, MixVPR mixvpr achieves improved performance, espeically in challenging Nordland dataset with varying weather and seasons. More recently, pre-trained features from visual foundation model (such as DINOv2 oquab2023dinov2 ) are concluded to be an excellent alternate for GF-VPR, and enable AnyLoc keetha2023anyloc achieves impressive retrieval performance in very different scene. Although GF-VPR methods can fully retrieve reference images, their retrieval accuracy is lower than SOTA LF-VPR methods as shown in Tab. 3. Adding verification with local features yue2 or jointly utilizing local feature and global feature fild2 can achieve pleased accuracy while retrieve most of reference images for each query image.
I2P-VPR: As for cross-modal I2P-VPR between query image and reference point cloud clip, a SOTA performance is achieved by I2P-Rec Li2023I2PRecRI as shown in Fig. 17. With the assistance of mature monocular depth estimation, I2P-Rec Li2023I2PRecRI transform query image and point cloud clip from scene map to BEV images, simplifying the cross-modal retrieval problem and leading to pleased performance. It also concludes such a BEV representation of point cloud can work better than point cloud representation, and better monocular depth estimation leads to better retrieval performance
@@ 8.3.1 MRL  for  IPA
Table 2: Recall over top-K candidates of different GF-VPR techniques on popular benchmarks.
Method dim Pitts250k-test pittsburgh MSLS-val MSLS SPED SPED Nordland nordland
@@
AVG netvlad ${}^{\dagger}$ 2048 62.6 82.7 88.4 59.3 71.9 75.5 54.7 72.5 77.1 4.4 8.4 10.4
GeM GeM ${}^{\dagger}$ 2048 72.3 87.2 91.4 65.1 76.8 81.4 55.0 70.2 76.1 7.4 13.5 16.6
NetVLAD netvlad ${}^{\dagger}$ 32768 86.0 93.2 95.1 59.5 70.4 74.7 71.0 87.1 90.4 4.1 6.6 8.2
AVG netvlad ${}^{\star}$ 2048 78.3 89.8 92.6 73.5 83.9 85.8 58.8 77.3 82.7 15.3 27.4 33.9
GeM GeM ${}^{\star}$ 2048 82.9 92.1 94.3 76.5 85.7 88.2 64.6 79.4 83.5 20.8 33.3 40.0
NetVLAD netvlad ${}^{\star}$ 32768 90.5 96.2 97.4 82.6 89.6 92.0 78.7 88.3 91.4 32.6 47.1 53.3
SPE-NetVLAD SPE-NetVLAD ${}^{\star}$ 163840 89.2 95.3 97.0 78.2 86.8 88.8 73.1 85.5 88.7 25.5 40.1 46.1
Gated NetVLAD vlaad ${}^{\star}$ 32768 89.7 95.9 97.1 82.0 88.9 91.4 75.6 87.1 90.8 34.4 50.4 57.7
CosPlace cosplace ${}^{\star}$ 2048 91.5 96.9 97.9 84.5 90.1 91.8 75.3 85.9 88.6 34.4 49.9 56.5
MixVPR mixvpr ${}^{\star}$ 2048 94.1 98.2 98.8 87.0 92.7 94.2 84.7 92.1 94.4 57.9 73.8 79.0
MixVPR mixvpr ${}^{\star}$ 4096 94.6 98.3 99.0 88.0 92.7 94.6 85.2 92.1 94.6 58.4 74.6 80.0
Method Baidu Mall BaiduMail Gardens Point GardenPoint2 17 Places 17places Pitts-30k pittsburgh St. Lucia RatSLAM Oxford oxford
@@
NetVLAD netvlad 53.10 70.51 58.50 85.00 61.58 77.83 86.08 92.66 57.92 72.95 52.88 74.87
CosPlace cosplace 41.62 55.02 74.00 94.50 61.08 76.11 90.45 95.66 99.59 99.93 91.10 99.48
MixVPR mixvpr 64.44 80.28 91.50 96.00 63.79 78.82 91.52 95.47 99.66 100 90.05 98.95
AnyLoc-GeM-DINOv2 keetha2023anyloc 50.13 70.55 88.00 97.50 63.55 79.56 77.04 87.28 76.91 89.34 81.15 97.38
AnyLoc-VLAD-DINO keetha2023anyloc 61.17 78.32 95.00 98.50 63.79 78.82 83.45 91.99 88.46 94.88 78.53 96.34
AnyLoc-VLAD-DINOv2 keetha2023anyloc 75.22 87.57 95.50 99.50 65.02 80.54 87.66 94.69 96.17 98.84 98.95 100
  • 1
    The results are cited from MixVPR mixvpr and AnyLoc keetha2023anyloc .
  • $\dagger$
    The results reported by the original publications.
  • $\star$
    The variants are trained on GSVcities dataset GSVCities using the same backbone network (ResNet-50 resnet ) with MixVPR mixvpr .
Methodology From Oxford fabmap KITTI KITTI Malaga malaga St. Lucia RatSLAM
@@
NetVLAD netvlad TPAMI’18 0.7148 0.4369 0.9315 0.9108 0.9632 0.3236 0.8046 0.7967
ResNet50-AP-GeM GeM TPAMI’19 0.9056 0.8485 0.9124 0.9470 0.9779 0.6011 0.7926 0.6936
Kazmi et al. Kazmi2019TRO TRO’19 0.7558 0.5109 0.9039 0.8141 0.9739 0.5098 0.8006 0.7255
@@
FABMAP 2.0 fabmap2 IJRR’11 0.3526 0.6038 0.8222 0.3712 0.6347 n.a. n.a. n.a.
DLoopDetector dbow-tro TRO’12 0.3059 n.a. 0.7243 0.5197 0.8971 0.7475 0.3722 0.3136
Tsintotas et al. AVWs ICRA‘18 0.6595 0.2988 0.9318 0.9420 0.9669 0.8799 0.2627 0.1507
iBoW-LCD ibow-lcd RAL’18 0.8825 0.7940 0.7650 0.5307 0.9553 0.5098 0.7002 0.8750
Tsintotas et al. BoTW RAL’19 n.a. n.a. 0.9750 0.9260 0.8897 0.8500 n.a. n.a.
Yue et al. yue1 IROS’19 0.8984 0.9229 0.9416 0.9494 0.9963 0.8605 n.a. n.a.
Han et al. lcd-pl JFR’21 0.3345 0.5982 0.9145 0.8651 n.a. 0.7251 n.a.
SLCD slcd TII’21 0.4088 0.7529 0.9753 0.8972 0.9313 n.a. n.a. n.a.
Yue et al. yue2 JFR’22 0.9127 0.9463 0.9569 0.9518 0.9963 0.8992 n.a. n.a.
@@
HTMap htmap TRO’17 0.7968 0.7360 0.9024 0.7588 0.9703 n.a. n.a. n.a.
FILD fild IROS’19 0.6648 0.7674 0.9123 0.8515 0.9338 0.5609 0.7606 0.6696
FILD++ fild2 JFR’22 0.9091 n.a. 0.9492 0.9542 0.9816 0.6274 0.8339 0.8136
Liu et al.liu ICRA’21 0.8601 0.9121 0.9302 0.9253 n.a. n.a. n.a. n.a.
@@
SeqSLAM seqslam ICRA’12 0.3186 0.5757 0.8758 0.1823 0.6068 n.a. n.a. n.a.
Bampis et al. Bampis2018IJRR IJRR’18 0.4963 0.8087 0.8947 0.8771 0.8015 0.3393 0.6093 0.4979
@@
    Most of the results are cited from recent papers yue2 and fild2 .

### v3

@@ VII-C Neural Radiance Field
NeRF implicitly represents a scene as a “radiance field”: a volumetric density that models the geometry of the scene, and a view-dependent color that models the appearance of occupied regions in the scene [269]. A Multiple Layer Perception (MLP) is usually utilized to takes the 3D position of a point P and the unit-norm viewing direction of that point v as inputs, and estimate the density $\sigma$ and color $c$ of the point, $(\sigma,c)\leftarrow\mbox{NeRF}(\textbf{P},\textbf{v})$. Scene map is implicitly represented by the parameters $\Theta$ of $\mbox{NeRF}(\cdot,\cdot)$. Given some RGB images $\{\mathcal{I}_{t}\}$ with known camera poses $\{\textbf{x}^{\mathcal{G}}_{t}\}$, NeRF is trained by minimizing a photometric loss $\mathcal{L}$ [269]:
@@
NeRF learns to synthesize novel views associated with any camera poses, enabling virtual view rendering and scene geometry recovery, thus have been applied in MRL in recent years. Some works directly uses pre-trained NeRF to estimate pose while other works utilize NeRF as a novel data augmentation method to enrich the training data.
NeRF as pose estimator: As the first framework to estimate pose using NeRF, iNeRF [270] directly optimizes 6 DoF poses with freezed pre-trained network parameters $\Theta$ by minimizing the photometric error between rendered pixels and observed pixels, which is an inverse problem of original NeRF [269]:
@@
In a very similar way, NeRF– [271] simplifies the training process of NeRF on forward-facing scenes by removing the need of camera pose and intrinsics. Later, SiNeRF [272] improves by leverage sinusoidal activations for radiance mapping and proposing a novel mixed region sampling to select ray batch efficiently. These joint pose estimation and NeRF training methods then motivate pose-free NeRF [273, 274].
@@
Many methods aiming at pose estimation only are also developed. In LATITUDE [6], authors trained a pose regressor through images generated from pre-trained NeRFs, which provides an initial pose guess for hierarchical localization without the requirement of saving reference images. In fine pose optimization stage, LATITUDE [6] minimizes the photometric error between the observed image and rendered image by optimizing the pose on the tangent plane. In Loc-NeRF [7], a pre-trained NeRF model is integrated into a Monte Carlo localization to measure weights to particles by image similarity between query image and rendered image. In NeRF-Loc [275], a NeRF conditioned on reference images takes 3D point as input and generates corresponding 3D descriptors so that 2D-3D matches can be obtained by matching. Precise 6 DoF pose is then estimated by a PnP solver.
@@
Figure 15: Instances of APR method enhanced by NeRF, a) Direct-PoseNet [276] using photometric consistency between query image and rendered image, while b) DFNet [277] using feature consistency.
NeRF as data augmentation: As a more straightforward way to apply NeRF in MRL, NeRF can be used to render virtual images at unseen viewpoints so that the overfitting problem of APR methods can be effectively solved and thus improves performances of APR [278, 276, 277]. LENS [278] applies pre-trained NeRF-W [279] as an offline data augmentation pipeline to enhance APR model. As an online augmentation way, Direct-PoseNet [276] minimizes the photometric errors between query image and virtual image rendered at estimated pose as shown in Fig. 15 a) while DFNet [277] uses feature map errors [277] as shown in Fig. 15 b), making APR method more precise and scalable. Enhanced by NeRF, APR methods can achieve significantly improved localization accuracy.
Summary: As a newly emerged technique, NeRF has two favourable characterises: 1) the implicit representation of scene; 2) the ability to render synthesis image at novel viewpoint. These two characterises present great potential in MRL researches. But we have to say the utilization of NeRF in MRL is still an opening question to be explored.
VIII Benchmarking Monocular Re-Localization
@@
In this section, the common evaluation metrics and datasets are introduced, and then we provide the performances of some SOTA MRL methods with various type of scene map on these benchmark. Looking back to all kind of MRL methods mentioned above, some of MRL methods retrieve reference image as query pose approximation while others perform precise pose estimation for query image. We name these two kind of MRL methods as Coarse Pose Approximation (CPA) and Fine Pose Estimation (FPE), respectively, as listed in Tab. I. These two kinds of MRL are generally tested using different evaluation metric. We provide introduction and evaluation respectively.
TABLE I: The categorisation of various MRL methods based on the function and their evaluation metrics.
Function
Method Evaluation Metric
Coarse Pose Approximation (CPA)
VPR (Sec. III-A), I2P-VPR (Sec. V-B) Precision (P), Recall (R), R@K, maxP@R=1
Fine Pose Estimation (FPE)
RPR (Sec. III-B), VL-MRL (Sec. IV), G-I2P (Sec. V-A), I2P-RPR (Sec. V-B), I2P-MLoc (Sec. V-B), HD-MRL (Sec. VI), NN-MRL (Sec. VII) Position error $\mbox{err}_{pos}$, Orientation error $\mbox{err}_{rot}$
TABLE II: Some datasets used as evaluation benchmarks in this paper.
Datasets Description Evaluated Methods
GP [280] campus, 3 traverses with the same trajectory (2 at daytime while 1 at nighttime) VPR (Sec. III-A)
SPEDTest [281] collected from all over the world, weather, seasonal and illumination changes
Nordland [14] natural and outdoor scenes, seasonal changes across different traverses
Pittsburgh [282] collected from Google Street View, significant viewpoint changes
MSLS [283] collected by ground vehicles, viewpoint and illumination changes
Oxford [79] including CityCentre and NewCollege, campus, fish-eye binocular images
St. Lucia [284] suburb, viewpoint and day-night appearance changes, dynamic objects
17-Places [285] indoor scenes, viewpoint changes
Baidu Mail [286] indoor scenes, reflectance and repetitive structures
Aachen Day-Night [33] query images taken at daytime and reference images taken at nighttime, significant illumination and viewpoint changes VL-MRL (Sec. IV)
InLoc [35] indoor scenes, large viewpoint and illumination changes, moving furniture, occlusions
KAIST [287] highly complex urban for AD HD-MRL (Sec. VI)
7Scenes [27] RGB-D images with GT poses are provided, 7 different indoor scenes ALL
Cambridge Landmarks [27] 6 medium-scale outdoor scenes
KITTI [288] collected by ground vehicles, 11 of 22 sequences have GT trajectories, both images and synchronized LiDAR scans are provided
VIII-A Evaluation metrics
VIII-A1 MRL  for  CPA
@@
Those MRL methods for CPA utilize evaluation metric as image retrieval task. A retrieved reference image/point cloud clip is defined as true positive if it was geometrically near the query images. The ground-truth is generally labeled by geometric distance based on GPS signals [200] or manual annotation [87, 88]. With the ground-truth, we can measure Precision (P) and Recall (R) metric, where precision denotes the ratio between true positives and all the retrieved references and recall is the ratio between the number of true positives and all the potentially matched query-references existed in this scene.
In some MRL application need to retrieve comprehensively, we measure the ratio of retrieved true positives over top-K candidates, R@K. In some applications that a mismatch will make the whole system dramatically fails, e.g., loop closure detection in SLAM [2], we turn to measure the maximum recall metric with 100% precision, maxP@R=1. The higher these two metrics, the better a MRL method performs.
VIII-A2 MRL  for  FPE
@@
Direct measurement about estimated poses is applied in those MRL methods aiming at FPE. The common evaluation metrics are position error $\mbox{err}_{pos}$ for translation $\textbf{t}^{\mathcal{G}}_{v}$ (in distance) and orientation error $\mbox{err}_{rot}$ for rotation $\textbf{R}^{\mathcal{G}}_{v}$ (in degree) between estimated results and ground-truth poses. The lower these two errors, the better a MRL method performs. In common scenes that vehicles move in 6 DoF, the position error and orientation error are both 3 DoF. But in AD scene for ground vehicles, we measure 2 DoF position error for lateral and longitudinal position, and 1 DoF orientation error for heading angle.
@@
VIII-B Dataset
@@
There are numerous datasets available for the evaluation of MRL methods and it is difficult to give a complete overview. Here we only introduce some datasets that we used to evaluate MRL methods in this survey, as listed in Tab. II.
VIII-B1 MRL  for  CPA
@@
The dataset used to evaluate VPR is required to include query and reference images that can be paired by geometric distance. So the dataset should have traversed the same scene multiple times, or the trajectory of the dataset should be “looped”, such as Gardens Point (GP) dataset [280], SPEDTest dataset [281], Nordland dataset [289], Pittsburgh dataset [282], Mapillary Street-Level Sequences (MSLS) dataset [283], Oxford CityCentre and NewCollege dataset [79], St. Lucia dataset [284], 17-Places dataset [285], Baidu Mail dataset [286], Tokyo 24/7 dataset [29], Synthia dataset [290], etc.
@@
To evaluate MRL for CPA (including both VPR and I2P-VPR), it requires the dataset to have both images and LiDAR sweeps. Since I2P-VPR is a newly emerged research, there is very few dataset proposed for the evaluation in this field. Researches in this field often use KITTI dataset [288]. Based on the given pose of each image in KITTI [288], ground-truth matches between query and reference image/point cloud can be labelled, and both VPR and I2P-VPR can be evaluated since synchronized LiDAR scans are also provided. We generally offline build LiDAR point cloud map using LiDAR SLAM to generate reference point cloud clip.
VIII-B2 MRL  for  FPE
Some datasets provide ground-truth pose for each image so that we can evaluate the MRL methods aiming at FPE. For example, the outdoor Aachen Day-Night dataset [33] and the indoor InLoc dataset [35] are usually utilized to evaluate VL-MRL methods. KAIST Urban dataset [287] provides LiDAR data and stereo image with various sensors targeting a highly complex urban environment. Some works [8, 226] build HD map for KAIST [287] so that it can be used to evaluate HD-MRL methods. KITTI dataset [288] provides ground-truth pose for each image, so it can also be used to evaluate MRL for FPE. Since multiple sensor data is provided, it can support evaluation for all kind of MRL. The indoor 7Scenes dataset [27] and outdoor Cambridge Landmarks dataset [27] provides ground-truth camera pose and 3D scene model, so they also support the evaluation of almost all the kinds of MRL methods. We compare most MRL methods on these two dataset to analyze the strengths and drawbacks of each kind of MRL methods. Except to these datasets, CMU Seasons and Extended CMU Seasons [33], RobotCar Seasons [33], ETH-Microsoft Dataset [291] are also widely used in evaluation, Map-free Localization dataset [21] provides a specialised benchmark for evaluating RPR methods when only one reference image is included in the scene map.
@@ VIII-B2 MRL  for  FPE
TABLE III: Recall over top-K candidates of different GF-VPR techniques on popular benchmarks.
Method dim Pitts250k-test [282] MSLS-val [283] SPED [281] Nordland [289]
@@
GeM [63] ${}^{\dagger}$ 2048 72.3 87.2 91.4 65.1 76.8 81.4 55.0 70.2 76.1 7.4 13.5 16.6
NetVLAD [19] ${}^{\dagger}$ 32768 86.0 93.2 95.1 59.5 70.4 74.7 71.0 87.1 90.4 4.1 6.6 8.2
GeM [63] ${}^{\star}$ 2048 82.9 92.1 94.3 76.5 85.7 88.2 64.6 79.4 83.5 20.8 33.3 40.0
NetVLAD [19] ${}^{\star}$ 32768 90.5 96.2 97.4 82.6 89.6 92.0 78.7 88.3 91.4 32.6 47.1 53.3
CosPlace [292] ${}^{\star}$ 2048 91.5 96.9 97.9 84.5 90.1 91.8 75.3 85.9 88.6 34.4 49.9 56.5
MixVPR [66] ${}^{\star}$ 4096 94.6 98.3 99.0 88.0 92.7 94.6 85.2 92.1 94.6 58.4 74.6 80.0
Method Baidu Mall [286] Gardens Point [280] 17 Places [285] Pitts-30k [282] St. Lucia [284] Oxford [293]
@@
NetVLAD [19] 53.10 70.51 58.50 85.00 61.58 77.83 86.08 92.66 57.92 72.95 52.88 74.87
CosPlace [292] 41.62 55.02 74.00 94.50 61.08 76.11 90.45 95.66 99.59 99.93 91.10 99.48
MixVPR [66] 64.44 80.28 91.50 96.00 63.79 78.82 91.52 95.47 99.66 100 90.05 98.95
AnyLoc-VLAD-DINOv2 [42] 75.22 87.57 95.50 99.50 65.02 80.54 87.66 94.69 96.17 98.84 98.95 100
[…]
Methodology From Oxford [79] KITTI [288] Malaga [31] St. Lucia [284]
@@
NetVLAD [19] TPAMI’18 0.7148 0.4369 0.9315 0.9108 0.9632 0.3236 0.8046 0.7967
ResNet50-AP-GeM [63] TPAMI’19 0.9056 0.8485 0.9124 0.9470 0.9779 0.6011 0.7926 0.6936
Kazmi et al. [296] TRO’19 0.7558 0.5109 0.9039 0.8141 0.9739 0.5098 0.8006 0.7255
@@
FABMAP 2.0 [80] IJRR’11 0.3526 0.6038 0.8222 0.3712 0.6347 n.a. n.a. n.a.
DLoopDetector [77] TRO’12 0.3059 n.a. 0.7243 0.5197 0.8971 0.7475 0.3722 0.3136
Tsintotas et al. [84] ICRA‘18 0.6595 0.2988 0.9318 0.9420 0.9669 0.8799 0.2627 0.1507
iBoW-LCD [82] RAL’18 0.8825 0.7940 0.7650 0.5307 0.9553 0.5098 0.7002 0.8750
SLCD [34] TII’21 0.4088 0.7529 0.9753 0.8972 0.9313 n.a. n.a. n.a.
Yue et al. [88] JFR’22 0.9127 0.9463 0.9569 0.9518 0.9963 0.8992 n.a. n.a.
@@
HTMap [86] TRO’17 0.7968 0.7360 0.9024 0.7588 0.9703 n.a. n.a. n.a.
FILD [297] IROS’19 0.6648 0.7674 0.9123 0.8515 0.9338 0.5609 0.7606 0.6696
FILD++ [74] JFR’22 0.9091 n.a. 0.9492 0.9542 0.9816 0.6274 0.8339 0.8136
Liu et al.[298] ICRA’21 0.8601 0.9121 0.9302 0.9253 n.a. n.a. n.a. n.a.
@@
SeqSLAM [14] ICRA’12 0.3186 0.5757 0.8758 0.1823 0.6068 n.a. n.a. n.a.
Bampis et al. [299] IJRR’18 0.4963 0.8087 0.8947 0.8771 0.8015 0.3393 0.6093 0.4979
@@
    Most of the results are cited from recent papers [88] and [74].